how many liters of 35% alcohol and 60% alcohol must be mixed together to make 40 liters of a solution that is 53% alcohol?
Answer by stanbon(75887)

how many liters of 35% alcohol and 60% alcohol must be mixed together to make 40 liters of a solution that is 53% alcohol?
------------------
Equaton::
alch + alch = alch
0.35x + 0.60(40-x) = 0.52(40)
------
Multiply thru by 100 to get:
35x + 60*40 - 0.60x = 52*40
-25x = -8*40
-----
x = (8/5)8
x = 64/5 = 12 4/5 liters (amt of 35% solution needed)
=====================
Cheers,
Stan H.
=====================